The following data for a hypothetical country in millions of dollars for the year 2020. Depreciation =200, Exports =150, public transfer =200, Gross domestic private investment =300, corporate income tax =100, factor receipts from abroad =400, gevenment expenditure on goods and services =250, interest income =800, compensation of employees =2600, net interest on government debit =50, indirect business taxes =100, factor payments to abroad =200, imports =200, proprietors incom(profits) =700, retained corporate profit =200, personal consumption expenditure =4500, personal taxes =100, social security contribution =50, rental income =600. A. Calculate the gross domestic product of the country I. Using the expenditure approach II. Using the income approach B, Calculate the gross nationalproduct of the country. E, calculate the rate the inflation rate of the economy if the consumers price index for 2019 is 100. F, find personal disposable income.


1
Expert's answer
2021-09-02T10:47:14-0400

a)

i)GDP @MP calculation by expenditure method (2020)



Consumption Expenditure = Personal consumption expenditure+ Personal taxes+ Contribution to social security

                        "= 4500+ 100+ 250 = 4850"

Investment Expenditure   = 300

Government Expenditure  = Govt. Expenditure - Interest on Government Debit "= 250- 50 = 200"

Net export "= Export - Import = 150-200 = -50."


ii) GDP@MP calculation by income method (2020)



Profit = Proprietors Income + Retain Corporate Profit + Corporate Income Tax = 700+200+100 = 1000 


b)

  • Gross National Product at market price of the country = GDP @MP of the country + Factors receipts from abroad- Factors receipts to abroad

                                                 "= 5300 + 400 - 200\\

 = \$ 5500 Million"



c)

"inflation \space =\frac{current cpi-past cpi}{past\space cpi}\times 100"

"=\frac{111-100}{100}\times 100=11"


d) "Personal\space Disposable\space Income =Personal\space Income − Personal\space Income\space Tax"


From the given information, 

Personal Income includes "=Rental \space Income + Proprietor's\space Income + Social\space Security \space Contribution + Interest \space Income + Compensation\space of \space employees" . 

Personal Income "= 600+700+50+800+2600"

Personal Income "= 4750"


Personal Income tax includes personal taxes and indirect business tax.

Personal Income Tax "=100+100"

Personal Income Tax "= 200"

 

Thus,Personal Disposable Income "= 4750-200"

Personal Disposable Income"= 4450"